# Teleportation=Translation: Continuous recovery of black hole information

### v2

@@ 3.3 Path B: The Canonical Haagerup–Kosaki Interpolation
(Note: Formally, the interpolated weight $\phi_{s}$ is constructed via the Connes-Takesaki Radon-Nikodym cocycle $D\phi_{1}:D\phi_{0}]_{t}$, and $\tilde{E}_{s}$ is derived from the resulting $L^{p}$ space structure.)
@@
These properties follow from the construction of the interpolation path where $\phi_{1}$ is a restriction of $\phi_{0}$ via a conditional expectation ($\phi_{1}=\phi_{0}\circ\tilde{E}$) Kosaki84 .
(1) Boundary Conditions: At $s=0$, the interpolation trivializes to the identity map on $\tilde{\mathcal{M}}$, implying $\tilde{E}_{0}=\text{id}$ and thus $\tilde{\mathcal{N}}(0)=\tilde{\mathcal{M}}$. At $s=1$, the construction recovers the original conditional expectation $\tilde{E}_{1}=\tilde{E}$, whose range is by definition the subalgebra $\tilde{\mathcal{N}}$.
(2) Nesting: Since $\phi_{1}$ is derived from $\phi_{0}$ via the conditional expectation $\tilde{E}$, the interpolation parameter $s$ acts as a strictly monotonic ordering parameter. The theory of non-commutative $L^{p}$ interpolation guarantees that the domains of the associated modular operators (and hence the ranges of the conditional expectations) form a nested family of subspaces, strictly decreasing as $s$ increases from 0 to 1.
(3) Idempotency: This is a direct consequence of the nesting property. Since $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ for $s^{\prime}\geq s$, the operator $\tilde{E}_{s^{\prime}}$ is a projection onto a subspace of the range of $\tilde{E}_{s}$. Therefore, projecting onto the larger space first ($\tilde{E}_{s}$) and then onto the smaller space ($\tilde{E}_{s^{\prime}}$) is equivalent to projecting directly onto the smaller space: $\tilde{E}_{s^{\prime}}\tilde{E}_{s}=\tilde{E}_{s^{\prime}}$. ∎
The canonical $L^{p}$-interpolation provides more than mathematical smoothness; it imposes a strict ordering on information. As proven in Theorem 3.6, the algebras satisfy the nesting property $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ for $s^{\prime}\geq s$. This monotonicity identifies $s$ as a coarse-graining scale: increasing $s$ corresponds to systematically discarding information (reducing the algebra), strictly analogous to the flow of a renormalization group. Conversely, the unitary operator $\tilde{U}(s)$ generated by this flow acts as an information recovery process, reversing this coarse-graining by translating information from the finer-grained algebra back into the decodable sector. This construction provides a mathematically rigorous canonical path between the two algebras, ensuring the correct boundary conditions for the unitary path.
@@ 3.4 The Unitary Path and its Generator
The distinction between the pedagogical path (Sec. 3.2) and the physical path (Sec. 3.3) determines the strength of our claims.
  • Pedagogical Path: The naive CP-interpolation is not idempotent. One cannot define a global Tomita–Takesaki structure. At best, one can analyze correlator-level diagnostics (as in Sec. 4.4) under the assumption that its $s=0$ generator matches the physical one.
  • Physical Path: The semifinite-envelope path $\tilde{E}_{s}$ is idempotent (Theorem 3.6). Its image $\tilde{\mathcal{N}}(s)$ is a true von Neumann subalgebra. Hence, the Tomita–Takesaki machinery applies globally at each $s$. (For a simple demonstration of the unitary goal in a Type I model, see Appendix B.)
We can now rigorously define the modular conjugation $J_{\tilde{\mathcal{N}}(s)}$ and the strongly continuous unitary path:
@@
Note that $\tilde{U}(0)=J_{\tilde{\mathcal{M}}}J_{\tilde{\mathcal{M}}}=\mathbf{1}$ and $\tilde{U}(1)=J_{\tilde{\mathcal{M}}}J_{\tilde{\mathcal{N}}}=U_{\Gamma}$.
The path $\tilde{U}(s)$ does not generally satisfy the group property $\tilde{U}(s+t)=\tilde{U}(s)\tilde{U}(t)$. Therefore, Stone’s Theorem does not directly apply to define a generator $\tilde{G}$ such that $\tilde{U}(s)=e^{-is\tilde{G}}$. Instead, we rigorously define $\tilde{G}$ as the infinitesimal generator (tangent vector) at $s=0$. However, for this definition to be meaningful, we must ensure that the path is sufficiently smooth (differentiable) and that the resulting operator is self-adjoint. Let $\mathcal{D}_{\text{core}}$ be the dense set of vectors that are analytic for the modular flow $\sigma_{t}$.
@@
The modular covariance of $\tilde{E}$ is the key technical property that guarantees the success of this construction. It ensures that analytic elements (those with entire modular flows) are mapped to analytic elements. Concretely, if $x\in\tilde{\mathcal{M}}$ has an entire extension $t\mapsto\sigma_{t}^{\phi_{0}}(x)$, then $\tilde{E}(x)$ also has an entire extension. This is why the common analytic core $\mathcal{D}_{\text{core}}$ is invariant. This invariance forms a valid dense core for $\tilde{G}$, which is essential for its definition and for the subsequent analysis of its properties in Sec. 4.
The immediate physical consequence of this core preservation is that the constructed path is sufficiently smooth to define a generator. Specifically, the invariant core ensures regularity at the modular boundary:
@@
The map $s\mapsto\tilde{U}(s)$ constructed via the canonical $L^{p}$-interpolation is real-analytic for $s\in(0,1)$. Moreover, for vectors in the common modular core $\mathcal{D}_{\text{core}}$, it is strongly differentiable at the boundary $s=0$.
@@
The proof relies on the analytic structure of the Haagerup–Kosaki interpolation and the properties of the modular core.
@@
    Analytic Extension: The interpolated spaces $L^{p}(\tilde{\mathcal{M}},\phi_{s})$ are defined via complex interpolation in the strip $0<\text{Re}(z)<1$. Consequently, the associated modular operators and conjugations $J_{s}$ depend analytically on the parameter $s$ within the open interval. This implies that for any vector $\xi\in\mathcal{D}_{\text{core}}$, the vector-valued function $s\mapsto\tilde{U}(s)\xi$ admits an analytic extension in $(0,1)$.
@@
    Core Preservation: As discussed in Remark 3.7, the modular covariance of the conditional expectations implies that the modular automorphism group of the subalgebra $\tilde{\mathcal{N}}(s)$ is simply the restriction of the global modular group. Consequently, the modular operators $\Delta_{s}$ share the same domain of entire analytic vectors as that associated with the ambient algebra. Thus, the actions of $J_{\tilde{\mathcal{N}}(s)}$ (and hence $\tilde{U}(s)$) preserve the common analytic core $\mathcal{D}_{\text{core}}$.
  3. 3.
    Differentiability at the Boundary: The differentiability at $s=0$ is guaranteed by the definition of the analytic core itself. Vectors $\xi\in\mathcal{D}_{\text{core}}$ are those for which the modular flow $t\mapsto\sigma_{t}^{\phi_{0}}(\xi)$ extends to an entire analytic function. Since the interpolation parameter $s$ corresponds to analytic continuation in the modular time (effectively a rotation $t\to-is$), the map $s\mapsto\tilde{U}(s)\xi$ is differentiable at $s=0$ strictly for $\xi\in\mathcal{D}_{\text{core}}$.
@@
Based on Lemma 3.8, we define the operator $\tilde{G}$ on the common modular core $\mathcal{D}_{\text{core}}$ as the derivative at the origin:
@@
Finally, we must prove that this operator is physically valid (i.e., self-adjoint).
@@
We rely on Nelson’s analytic vector theorem and the analyticity established in Lemma 3.8.
@@
    Symmetry: Since $\tilde{U}(s)$ is unitary for real $s$, differentiating the identity $\tilde{U}(s)^{*}\tilde{U}(s)=\mathbf{1}$ at $s=0$ yields $\tilde{U}^{\prime}(0)^{*}+\tilde{U}^{\prime}(0)=0$, implying $\tilde{U}^{\prime}(0)^{*}=-\tilde{U}^{\prime}(0)$. Thus, $\tilde{G}=i\tilde{U}^{\prime}(0)$ is a symmetric operator ($\tilde{G}^{*}=\tilde{G}$) on $\mathcal{D}_{\text{core}}$.
@@
    Analytic Vectors: From Lemma 3.8, for any vector $\xi\in\mathcal{D}_{\text{core}}$, the function $s\mapsto\tilde{U}(s)\xi$ is analytic at $s=0$. Since $\tilde{U}(s)$ is locally generated by $\tilde{G}$, the power series expansion of this function guarantees that $\xi$ is an analytic vector for the operator $\tilde{G}$. Thus, $\mathcal{D}_{\text{core}}$ constitutes a dense set of analytic vectors.
@@
    Conclusion: By Nelson’s analytic vector theorem Nelson59 , a symmetric operator possessing a dense set of analytic vectors is essentially self-adjoint. Therefore, the closure of $\tilde{G}$ is the unique self-adjoint generator.
@@
The canonical $L^{p}$-interpolation path employed here is rigorously defined without inherent ambiguity. The path is uniquely determined by the initial and final algebras and weights, making $\tilde{G}$ a canonical object of the theory, invariant under unitary conjugations that preserve the inclusion. This $\tilde{G}$ is the well-defined, self-adjoint generator of our continuous teleportation protocol.
4 ALGEBRAIC VERIFICATION: $\tilde{G}=2P$
The existence of the uniquely defined self-adjoint generator $\tilde{G}$ (established in Sec. 3.4) allows us to rigorously verify the core physical hypothesis: the identification of the teleportation generator with the modular momentum. This section provides a full functional-analytic proof of the identity $\tilde{G}=2P$ for the general Type III case. We structure the proof to bridge physical motivation with mathematical rigor. We begin by analyzing the half-sided modular inclusion (HSMI) setting to provide a robust motivation for the factor of 2 (Sec. 4.1). Subsequently, we present the rigorous derivation using modular perturbation theory, establishing the identity as a theorem for closed operators (Sec. 4.2). We further quantify the structural robustness of this result against perturbations using non-commutative geometry (Sec. 4.3). Finally, we propose a concrete correlation-function test applicable to holographic models (Sec. 4.4).
4.1 Verification in Half-Sided Modular Inclusions
A strong theoretical foundation for our conjecture arises from the theory of HSMI, which frequently appear in algebraic quantum field theory models (e.g., wedge algebras). In this highly symmetric setting, the connection between the canonical shift and modular operators is not merely a conjecture but a rigorous theorem, established by Borchers and Araki Araki73 ; Borchers92 ; Longo89 .

### v3

@@ 3.3 Path B: The Canonical Haagerup–Kosaki Interpolation
The intermediate expectations $\tilde{E}_{s}$ are uniquely determined by the analytic continuation of the Connes–Takesaki Radon–Nikodym cocycle $D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}$. Specifically, if $h=d\tilde{\phi}_{1}/d\tilde{\phi}_{0}$ denotes the non-commutative Radon–Nikodym derivative, the interpolation is driven by the relation $D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}=h^{it}$. This relation underpins the analyticity of the Haagerup $L^{p}$-interpolation path at the boundaries (particularly at $s=0$) and ensures that the interpolated weights $\tilde{\phi}_{s}$ are uniquely determined by the resulting non-commutative $L^{p}$ space structure.
The construction in Definition 3.5 has a clear physical interpretation in terms of non-commutative geometry. The Radon–Nikodym derivative relation implies that the interpolated weight $\tilde{\phi}_{s}$ satisfies the exact power-law relation regarding its density:
$\frac{d\tilde{\phi}_{s}}{d\tilde{\phi}_{0}}\;=\;h^{s}\;=\;\left(\frac{d\tilde{\phi}_{1}}{d\tilde{\phi}_{0}}\right)^{s}.$ (15)
Heuristically, this signifies that $\tilde{\phi}_{s}$ acts as a non-commutative geometric mean between the reference weight $\tilde{\phi}_{0}$ and the coarse-grained weight $\tilde{\phi}_{1}$. Although formal products of weights are not defined in the operator algebra, we may conceptually visualize this interpolation as:
$\tilde{\phi}_{s}\;\sim\;\tilde{\phi}_{1}^{\,s}\,\tilde{\phi}_{0}^{\,1-s}~.$ (16)
This geometric averaging property ensures that the path $\tilde{\mathcal{N}}(s)$ follows the geodesic of the underlying modular structure, minimizing the information distance between the full algebra and the subalgebra.
@@
These properties follow directly from the analytic structure of the Haagerup-Kosaki interpolation. Unlike naive spectral truncations, this path relies on complex interpolation between the weights $\tilde{\phi}_{0}$ and $\tilde{\phi}_{1}$, ensuring structural continuity via the modular automorphism group.
(1) Boundary Conditions: The interpolation relies on the analytic continuation of the Radon-Nikodym derivative $h$ defined above. At $s=0$, the interpolation yields the unperturbed reference weight $\tilde{\phi}_{0}$, implying that the associated projection is the identity map ($\tilde{E}_{0}=\text{id}_{\tilde{\mathcal{M}}}$). Conversely, at $s=1$, the construction converges in norm to the restricted weight $\tilde{\phi}_{1}$, thereby recovering the original conditional expectation $\tilde{E}_{1}=\tilde{E}$ and its range $\tilde{\mathcal{N}}$. This smooth limit behavior eliminates the discontinuity at $s=0$ that plagues spectral cut methods.
(2) Nesting: This property arises from the definition $\tilde{\phi}_{1}=\tilde{\phi}_{0}\circ\tilde{E}$. The operator $h$ essentially acts as a density operator representing the projection $\tilde{E}$. The interpolation corresponds to powers $h^{s}$, which serve as a soft filter whose intensity increases with $s$. Since $0\leq h\leq 1$ (in the appropriate operator sense), increasing $s$ strictly contracts the support of the associated expectations. Thus, for $s^{\prime}\geq s$, the range of $\tilde{E}_{s^{\prime}}$ is contained within the range of $\tilde{E}_{s}$.
(3) Idempotency: This follows directly from the nesting property. Since $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ for $s^{\prime}\geq s$, the operator $\tilde{E}_{s^{\prime}}$ projects onto a subspace contained within the range of $\tilde{E}_{s}$. Thus, applying the coarser filter $\tilde{E}_{s}$ before the finer filter $\tilde{E}_{s^{\prime}}$ leaves the latter invariant: $\tilde{E}_{s^{\prime}}\tilde{E}_{s}=\tilde{E}_{s^{\prime}}$. ∎
The canonical $L^{p}$-interpolation provides more than mathematical smoothness; it imposes a strict ordering on information. The nesting property $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ established in Theorem 3.6 identifies $s$ as a coarse-graining scale: increasing $s$ corresponds to systematically discarding information, strictly analogous to the flow of a renormalization group. Conversely, the unitary operator $\tilde{U}(s)$ generated by this flow acts as an information recovery process, reversing this coarse-graining by translating information from the finer-grained algebra back into the decodable sector. Thus, this construction not only connects the two algebras but also establishes the rigorous framework required for the unitary recovery path.
@@ 3.4 The Unitary Path and its Generator
The distinction between the pedagogical path (Sec. 3.2) and the physical path (Sec. 3.3) is not merely technical but conceptual. What is required is not just a continuous family of completely positive maps connecting the endpoints, but a one-parameter family of conditional expectations whose ranges remain von Neumann algebras. Only then do the modular objects exist along the entire path allowing for the rigorous definition and differentiation of the unitary flow that implements information recovery.
  • Pedagogical Path (Algebraic Limitation): The linear CP-interpolation $E_{s}=(1-s)\mathrm{id}+sE$ provides a mathematically well-defined path of maps. However, as illustrated in the $M_{2}(\mathbb{C})$ model (Appendix B), this linearity generally does not preserve the idempotency condition ($E_{s}^{2}\neq E_{s}$) for intermediate values of $s$. This property holds regardless of the algebra’s Type or the existence of a trace. As a result, the image of $E_{s}$ does not form a von Neumann subalgebra, which precludes the definition of the relative modular operator associated with the inclusion necessary to generate the unitary flow.
  • Physical Path (Algebraic Consistency): Conversely, the path $\tilde{E}_{s}$, defined by the canonical $L^{p}$-interpolation within the semifinite envelope, is idempotent by construction (Theorem 3.6). Since its image $\tilde{\mathcal{N}}(s)$ forms a von Neumann subalgebra for all $s\in[0,1]$, modular theory remains applicable throughout the interpolation. This structure ensures that the path represents a genuine flow of algebras rather than a simple interpolation of states, enabling the construction of the generating Hamiltonian even within the Type III setting.
The structural consistency of the physical path enables an explicit definition of the modular conjugation $J_{\tilde{\mathcal{N}}(s)}$ and the strongly continuous unitary path:
@@
This specific construction represents the continuous counterpart to the operational realization of the discrete information recovery process discussed in vdHV and summarized in Sec. 2.2. Applying the logic established in Step 3 of the protocol, the modular conjugation acts as a teleportation map at each $s$, relocating the instantaneous relative commutant $\tilde{\mathcal{N}}(s)^{\prime}\cap\tilde{\mathcal{M}}$ to the corresponding basic extension’s commutant $\tilde{\mathcal{M}}^{\prime}\cap\tilde{\mathcal{M}}_{1}(s)$, where
$\tilde{\mathcal{M}}_{1}(s):=\langle\tilde{\mathcal{M}},e_{\tilde{\mathcal{N}}(s)}\rangle.$ (18)
Consequently, the flow generated by $\tilde{U}(s)$ explicitly implements this transport: it progressively decodes the information hidden in the relative commutant sector, mapping it back into the decodable sector, thereby effectively reversing the coarse-graining induced by the inclusion. Crucially, the analytic nature of the Haagerup–Kosaki interpolation guarantees the boundary conditions $\tilde{U}(0)=\mathbf{1}$ and $\tilde{U}(1)=U_{\tilde{\Gamma}}$, ensuring that the path continuously connects the identity to the target canonical shift.
The path $\tilde{U}(s)$ does not generally satisfy the group property $\tilde{U}(s+t)=\tilde{U}(s)\tilde{U}(t)$. Consequently, Stone’s Theorem does not directly apply to define a generator $\tilde{G}$ such that $\tilde{U}(s)=e^{-is\tilde{G}}$. Instead, we define $\tilde{G}$ as the infinitesimal generator (the tangent vector) at $s=0$. For this operator to be well-defined and physically meaningful, we must ensure that the path is strongly differentiable on a suitable domain and $\tilde{G}$ is essentially self-adjoint. To establish these properties, we consider a dense set of vectors $\mathcal{D}_{\text{core}}\subset\mathcal{H}$ that are analytic with respect to the modular flow $\sigma_{t}^{\tilde{\phi}_{0}}$. This choice of core allows us to control the domain of the generator and provides the necessary foundation for proving its self-adjointness in Lemma 3.10.
@@
The modular covariance of the conditional expectation $\tilde{E}_{s}$ is the pivotal property ensuring the structural consistency of this construction. Specifically, the fact that $\tilde{E}_{s}$ commutes with the modular flow $\sigma_{t}^{\tilde{\phi}_{0}}$ guarantees that analyticity is preserved under projection: analytic elements—those possessing entire modular continuations—are mapped to analytic elements within the subalgebra. Concretely, if $x\in\tilde{\mathcal{M}}$ admits an entire extension $t\mapsto\sigma_{t}^{\tilde{\phi}_{0}}(x)$, its image $\tilde{E}_{s}(x)$ likewise admits an entire extension. Consequently, the common analytic core $\mathcal{D}_{\text{core}}$ remains invariant under the interpolated expectations for all $s$. This invariance provides the stable, dense domain required to strictly define $\tilde{G}$ as a derivation, serves as the foundation for proving its self-adjointness in Lemma 3.10, and facilitates the broader analysis of its properties in Sec. 4.
The immediate physical consequence of this core preservation is that the constructed path is sufficiently smooth to define the generator $\tilde{G}$. Specifically, the invariant core $\mathcal{D}_{\text{core}}$ ensures regularity with respect to the modular flow, allowing for well-defined differentiation at $s=0$:
@@
The map $s\mapsto\tilde{U}(s)$ constructed via the canonical $L^{p}$-interpolation is real-analytic for $s\in(0,1)$. Moreover, for vectors in the common modular core $\mathcal{D}_{\text{core}}$, this map is strongly differentiable at the boundary $s=0$.
@@
The proof relies on the analytic structure of the Haagerup–Kosaki interpolation and the properties of the modular domain.
@@
    Analytic Extension inside the Interval: The interpolated spaces $L^{p}(\tilde{\mathcal{M}},\tilde{\phi}_{s})$ are defined via complex interpolation in the strip $0<\text{Re}(z)<1$, where $z$ extends the real parameter $s$ into the complex plane. Consequently, the structural maps, including the modular conjugations $J_{\tilde{\mathcal{N}}(s)}$, depend analytically on the parameter $s$ within the open interval. This implies that for any vector $\xi\in\mathcal{D}_{\text{core}}$, the vector-valued function $s\mapsto\tilde{U}(s)\xi$ is real-analytic in $(0,1)$.
@@
    Differentiability at the Boundary: The differentiability at $s=0$ hinges on the invariance of the analytic core established in Remark 3.7. Recall that vectors $\xi\in\mathcal{D}_{\text{core}}$ are characterized by the property that their modular flow $t\mapsto\sigma_{t}^{\tilde{\phi}_{0}}(\xi)$ extends to an entire analytic function. Since the interpolation parameter $s$ corresponds to the analytic continuation of the modular group to the imaginary axis (mathematically relating $s$ to $it$), the existence of the derivative at $s=0$ is equivalent to the existence of the analytic continuation. The invariance of $\mathcal{D}_{\text{core}}$ ensures that this analytic structure is preserved along the path, guaranteeing that the limit defining the derivative exists strictly for vectors within this core.
@@
Motivated by the differentiability established in Lemma 3.8, we define the operator $\tilde{G}$ on the common modular core $\mathcal{D}_{\text{core}}$ as the tangent vector at the origin:
@@
Finally, we establish the physical validity of this operator.
@@
We invoke Nelson’s analytic vector theorem Nelson59 , utilizing the properties of the modular core $\mathcal{D}_{\text{core}}$.
@@
    Symmetry: Since $\tilde{U}(s)$ is unitary for real $s$, differentiating the identity $\tilde{U}(s)^{*}\tilde{U}(s)=\mathbf{1}$ at $s=0$ yields $\tilde{U}^{\prime}(0)^{*}\tilde{U}(0)+\tilde{U}(0)^{*}\tilde{U}^{\prime}(0)=0$. Using $\tilde{U}(0)=\mathbf{1}$, this implies $\tilde{U}^{\prime}(0)^{*}=-\tilde{U}^{\prime}(0)$. Thus, $\tilde{G}=i\tilde{U}^{\prime}(0)$ is a symmetric operator ($\tilde{G}^{*}=\tilde{G}$) on the dense domain $\mathcal{D}_{\text{core}}$.
@@
    Analytic Vectors: Recall that the generator $\tilde{G}$ arises from the analytic continuation of the Radon-Nikodym cocycle (effectively the logarithm of the relative modular operator $h$). The domain $\mathcal{D}_{\text{core}}$ is explicitly defined as the set of vectors that are analytic with respect to the modular flow generated by $\tilde{\phi}_{0}$ (and by extension, $h$). Mathematically, vectors that are analytic for a one-parameter group $h^{it}$ are analytic vectors for its generator $\log h$. Since $\tilde{G}$ is linearly related to this generator at $s=0$, the set $\mathcal{D}_{\text{core}}$ constitutes a dense set of analytic vectors for $\tilde{G}$.
@@
    Conclusion: By Nelson’s analytic vector theorem, a symmetric operator possessing a dense set of analytic vectors is essentially self-adjoint. Therefore, the closure of $\tilde{G}$ defines a unique self-adjoint quantum generator.
@@
The canonical $L^{p}$-interpolation path employed here is defined without ambiguity. It is uniquely determined by the initial reference weight and the target subalgebra, independent of arbitrary choices such as basis vectors or auxiliary cutoffs. Consequently, $\tilde{G}$ is a canonical object of the theory, invariant under unitary conjugations that preserve the inclusion structure. This establishes $\tilde{G}$ as the well-defined, self-adjoint generator of the continuous teleportation protocol.
4 DERIVATION OF THE OPERATOR IDENTITY $\tilde{G}=2P$
Having established the definition of the self-adjoint generator $\tilde{G}$ in Sec. 3.4, we now turn to the central objective of this work: establishing the identification of the teleportation generator with the modular momentum. This section presents a derivation of the identity $\tilde{G}=2P$ for general Type III algebras. We begin by analyzing the half-sided modular inclusion (HSMI) setting to motivate the specific form of the identity, particularly the factor of 2 (Sec. 4.1). Subsequently, we employ modular perturbation theory to derive the identity as an exact relation between closed operators (Sec. 4.2). We further examine the geometric stability of this result within the framework of non-commutative geometry (Sec. 4.3) and conclude by proposing a correlation function test applicable to holographic models (Sec. 4.4).
4.1 Motivation from Half-Sided Modular Inclusions
A strong theoretical foundation for the identification of the teleportation generator with modular momentum, as proposed by vdH-V vdHV , is found in highly symmetric settings, exemplified by HSMI. In this symmetric context, the connection between the canonical shift and modular operators is not merely a hypothesis but a rigorous theorem, established by Borchers and Wiesbrock Borchers92 ; Wiesbrock93 .